# Adv-BMT: Bidirectional Motion Transformer for Safety-Critical Traffic Scenario Generation

### v1

@@ 3.5 Adv-BMT
While existing works select a convenient neighbor agent and modify the behavior to attack the ego agent, Adv-BMT inserts new agents (ADV) with diverse collision initializations for different opponent interactions. We visualize an example result in 4. Formally, we define the collision state as to include the position, time, velocity, and heading at the collision step. The collision time is sampled from the first second to the last timestep of the ego trajectory length. Similarly, ADV’s collision headings are randomly sampled. ADV’s collision position can be calculated from collision heading and the ego vehicle’s collision position, with the ADV contour intersecting with the ego vehicle’s contour. Last but not least, collision velocity is also calculated from sampling a variable offset from ego vehicle’s speed at collision step.
Reverse Predictions
We use the sampled collision states as inputs of BMT for reverse predictions, and obtain ADV’s trajectory from predictions to augment the original scene with a new adv agent. During BMT inference, we teacher-force the ground-truth token sequence of ego agent and traffic agents, to keep the originality of the scenario as much as possible (since BMT could predict agents to driving on a different lane choice, which makes it a totally different scenario). By teacher-forcing traffic vehicle agents, BMT reconstructs the adversarial agent trajectory to fit in the input scenario. Reconstructing entirely new scenarios is out of the scope of our current work, but BMT is able to realize both normal and safety-critical generations through forward and reverse predictions.
Rule-based Rejection Samplings
@@
We design Adv-BMT to have diverse initializations, which do not guarantee the realism of the collision outcomes. To address this issue, we implement a rule-based rejection mechanism for ADV candidates. We first measure the driving distance and average speed of ADV candidates; if it moves too short and mostly wanders at the designated position waiting for the ego vehicle, then it is considered invalid. Meanwhile, we check the max curvature (the rate of change for heading): given a candidate ADV trajectory, we compute the curvature constraint using:$\kappa_{t}={\Delta\theta_{t}}/{\Delta s_{t}},$ where $\Delta\theta_{t}$ is the absolute heading change between time steps, and $\Delta s_{t}$ is the displacement between consecutive positions. Adv-BMT rejects a prediction if $\max_{t}(\kappa_{t})>\kappa_{\text{threshold}},$where $\kappa_{\text{threshold}}=0.8$ is a predefined curvature limit that we found useful. We check the curvature to ensure it does not exceed our defined threshold, and reject candidate ADVs with unrealistic changes in headings. With our straightforward rule-based rejection sampling mechanism, we are able to maintain only realistic collision events between the ADV and the ego vehicle.
Figure 3: BMT predicts both future and history agent trajectories.
Figure 4: Adv-BMT generates diverse collision scenarios.
Forward refinement for Reactive Traffic
In real-world collision events, neighbor vehicles would be affected by the collision vehicles and would have reactions like sudden lane changes and brakings. With this concern, we implement Bi-directional Adv-BMT, with a forward refinement to re-simulate traffic agents upon predicted ADV trajectory. In forward refinement, all traffic agents would be predicted and modified in the output scenario, with teacher-forcing motions of collision participants (ego and ADV agents). In experiments, we refer to it as Adv-BMT (Bi) to indicate bi-directions.
@@ 4 Experiments
In section 4.1, we evaluate BMT model on motion prediction performance. In section 4.2, we compare Adv-BMT with baseline methods. In section 4.3 we train a reinforcement learning agent in different safety-critical augmented datasets. For augmentations, we use ground-truth traffic data from the Waymo Open Motion Datasets (WOMD) ettinger2021largescaleinteractivemotion .
4.1 Realism of Adv-BMT
Settings
We measure BMT on four different settings: Forward prediction with or without agent history, Reverse prediction without agent history, and bi-directional predictions. We measure on open-loop scenario generation metrics with 6 prediction modes on each scenario: scenario final displacement errors (SFDE) which measures the average displacement error at the final step on the best prediction mode, and scenario average displacement errors (SADE) which measures the averages according to all time steps on the best mode. We also measure Final Displacement Diversity (FDD) and average displacement diversity (ADD) for prediction diversities across different prediction modes: FDD measures the spread of final positions and ADD measures trajectory variations over time. Finally, we measure the overall collision rates for realistic interactions. The results are presented in Table 1.
Table 1: Evaluation for BMT on scenario realism on different prediction modes.
Context SFDE $\downarrow$ SADE $\downarrow$ minSFDE $\downarrow$ minSADE $\downarrow$ FDD $\uparrow$ ADD $\uparrow$ Collision
Forward 1s 3.043 1.019 1.908 0.693 9.241 3.240 0.026
Forward 0s 3.180 1.297 2.144 0.927 10.263 4.177 0.028
Reverse 0s 3.411 1.380 2.158 0.975 8.234 2.913 0.031
Bidirections 0s 3.059 2.111 1.831 1.536 7.618 6.270 0.035
Analysis
Results from Table 1 indicate that BMT performs realistic scenario generations in both forward predictions and reverse predictions. Furthermore, by conditioning on a 1-second history, the forward prediction achieves better SADE and SFDE scores, indicating better temporal consistencies. Due to design of Adv-BMT with agent initializations at the last step, it is impossible for reverse predictions to access agent histories. Compared to single-pass prediction, bi-directional prediction exhibit lower prediction accuracy in ADE metrics, indicating a harm in the realism of generations.
@@ 4.1 Realism of Adv-BMT
In terms of prediction diversities, Forward prediction exhibits higher FDD and ADD scores than reverse prediction. This indicates that predicting histories is more constrained, but predicting future motions encourages more explorations for variant directions. This further validates the necessity of Adv-BMT’s diverse collision initialization design to force different collision outcomes. Despite these differences, the overall collision rate remains reasonable and comparable for all prediction modes, and indicates that both prediction modes effectively generate realistic interactions.
Figure 5: Comparison among adversarial behaviors of different methods.
4.2 Safety-critical Evaluations
In following experiments, We compare Adv-BMT with three baseline methods for safety-critical augmentations, namely STRIVE rempe2022generatingusefulaccidentpronedriving , CAT zhang2023catclosedloopadversarialtraining , and SEAL stoler2025sealsafeautonomousdriving . We evaluate the quality and efficiency of generated collisions among all methods. We measure the adversarial attack success rate (SDC-ADV collision rate), collision rates between adversary agents and traffic agents (ADV-BV Collision rate), indicating interaction realism. Additionally, we evaluate the difference between the predicted distribution and the ground-truth distribution on velocities (VEL JSD) and accelerations (ACC JSD). These two metrics also indicate the realism of generated adversarial agent motions.
Settings
We randomly select 500 WOMD scenarios, each as a base environment for adversarial augmentations. For each scenario, each method predicts six different adversarial trajectories. To be consistent with baseline methods in this evaluation only, we force Adv-BMT to choose an existing neighbor agent and modify its behavior as the predicted adversary.
Table 2: Evaluation for Safety-critical Scenarios
@@ 4.2 Safety-critical Evaluations
CAT zhang2023catclosedloopadversarialtraining 0.470 0.271 0.137 0.480 0.0 0.0
SEAL stoler2025sealsafeautonomousdriving 0.582 0.261 0.239 0.561 0.0 0.0
STRIVE rempe2022generatingusefulaccidentpronedriving 0.534 0.358 0.117 0.432 0.004 0.001
Adv-BMT 1.000 0.391 0.142 0.435 0 2.344
Adv-BMT (Bi) 0.997 0.373 0.188 0.534 0.038 2.257
Quantitative Results
Table 2 shows the quantitative results for scenario generation metrics of all methods. Two settings of Adv-BMT outperform baselines in terms of adversarial attack success rate. This validates Adv-BMT’s design of collision initialization + reverse predictions, which couldn’t be achieved by other methods. While Adv-BMT generates highly adversarial behaviors, it also preserves diversity of traffic interactions compared to baselines. Note tha FDD and SDD metrics suggest that our baselines generate nearly the same adversarial trajectory on the given scene, while Adv-BMT is able to generate diverse behaviors at different prediction mode. The JSD metrics suggests slightly better realism on velocities and accelerations for baselines compared to Adv-BMT, showing the trade-off between diversity and velocity/acceleration realism. BMT model is able to achieve realistic motion predictions indicted by Waymo Open Sim Agent Challenge metrics montali2023waymoopensimagents , which we append in our appendix. Overall, our safety-critical evaluations indicate that Adv-BMT is an efficient framework for realistic, diverse, and safety-critical generations.
Visualizations
We simulate Adv-BMT scenarios in MetaDrive li2022metadrivecomposingdiversedriving simulators using both BEV and 3D rendering. In Fig. 4, we show diverse Adv-BMT adversarial behaviors in the collision directions on the same ego agent in an intersection scenario from WOMD. Adversarial agents (in red) come from different lanes and collide with the ego agent (in green) at different angles and facets. Besides, the generated adversarial agent follows the traffic rules and maintains realistic driving patterns. Apart from diverse collision directions, Adv-BMT generates collisions at diverse timings. Adv-BMT adversarial agents cover collisions with the ego before entering the intersection, during the middle of the intersection, and after turning to the main lane. With a limited number of real-world maps and logs, Adv-BMT is able to generate diverse collision behaviors and timings, making it suitable for AD testing and adversarial training. A visual comparison among Adv-BMT and baseline methods is shown in Fig. 5, where other methods fail to generate an adversarial trajectory but Adv-BMT is able to insert a realistic ADV and triggers collisions .
Generation Speeds
Adv-BMT and CAT generate safety-critical scenarios significantly faster than other baselines. Adv-BMT generates one scenario in 1.02s compared to SEAL’s 2.36s and STRIVE’s 9.53s, achieving great efficiency gains. With details shown in Table 7 in Appendix.
@@ 4.3 Adversarial Learning
To validate the value of Adv-BMT scenarios in downstream autonomous driving (AD) tasks, we train a reinforcement learning agent within augmented scenarios with collision interactions generated by Adv-BMT and baseline methods. To determine the quality of augmented training scenarios, we measure both driving performance and safety performance of the learned AD agent compared to learning in original training set. In our experiment, we conduct two sets of trainings. (1) open-loop training: agent is trained in the fixed training set with adversarial scenarios generated based on ground-truth ego trajectories. (2) closed-loop training: an adaptive adversarial agent attacks on current ego agent based on its recent rollout trajectory records. The adaptive adversarial agent’s motion is generated by Adv-BMT and baseline. Results of both trainings are shown in table 3 and table 4 respectively.
Settings
The training set contains 500 real-world scenarios randomly selected from WOMD training set. Each waymo scenario is augmented with one collision scenario, so that the ratio between waymo and safety-critical scenes is 1:1. For each method, we augments the original dataset into a new training set. For the adaptive adversarial learning experiment, we implement an adaptive generator for Adv-BMT and CAT zhang2023catclosedloopadversarialtraining (discard other methods due to lower inference time). We train a Twin Delayed DDPG (TD3) agent for 1 million steps using 8 random seeds to ensure robustness. We conduct the training in the MetaDrive simulator and use the TD3 algorithm for reinforcement learning (hyper-parameters listed in Appendix). We measure the average reward, average cost, average route completion rate (Compl.), and average episode cost (cost sum) for driving performance measure. We further conducted ablation studies with results shown in Appendix.
Quantitative Results
To evaluate the impact of adversarial training using Adv-BMT-generated scenarios, we assess policy performance across two distinct validation environments: (1) 100 Waymo Validation Environments, which consist of unmodified real-world driving scenarios from WOMD validation set, and (2) 100 Adv-BMT Environments, which is the augmented collision scenarios from the 100 validation scenes. Table 3 presents the evaluation results, demonstrating how training with Adv-BMT improves policy resilience in both standard and adversarial settings. A detailed analysis of these results follows in the subsequent sections.
Table 3: Open-loop learning evaluation.
@@
Training Scenarios Reward $\uparrow$ Cost $\downarrow$ Compl. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
@@
Adv-BMT (Bi) 33.22 $\pm$ 1.83 0.36 $\pm$ 0.03 0.74 $\pm$ 0.03 0.12 $\pm$ 0.02 1.39 $\pm$ 0.22
(b) Evaluation in the Adv-BMT Environments
Training Scenarios Reward $\uparrow$ Cost $\downarrow$ Compl. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
@@
Adv-BMT (Bi) 39.55 $\pm$ 2.94 0.59 $\pm$ 0.04 0.65 $\pm$ 0.02 0.27 $\pm$ 0.04 2.74 $\pm$ 0.54
Table 4: Closed-loop learning results
@@
Adaptive Generator Reward $\uparrow$ Cost $\downarrow$ Comp. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
@@
Adaptive Generator Reward $\uparrow$ Cost $\downarrow$ Comp. $\uparrow$ Coll. $\downarrow$ Cost Sum $\downarrow$
@@
Analysis
1) Improved Policy Robustness in Adversarial Settings. Table 3 presents the evaluation results comparing policies trained with Adv-BMT-generated scenarios against baselines in both real-world and safety-critical validation environments. Adv-BMT demonstrates superior performance in reducing collision rates while maintaining high completion rates. In the collision validation environments, Adv-BMT-trained policies achieve the lowest collision rate while preserving competitive completion rates. This highlights the effectiveness of Adv-BMT in exposing policies to diverse and high-risk scenarios, leading to improved safety performance. Furthermore, Adv-BMT achieves the lowest sum cost, suggesting that policies trained on Adv-BMT-generated data are more efficient in handling safety-critical interactions. Besides, the results from adaptive generators shows an slight improvement in performance compared to training in fixed scenario sets.
In the open-loop learning, among the baselines, STRIVE achieves the highest reward (37.72) in collision environments, but this comes at the cost of a higher collision rate (0.29). SEAL and CAT exhibit higher costs and collision rates, which suggests that adversarial training leads to suboptimal behaviors. In contrast, agent training with Adv-BMT attacks results in improvement on both safety and task completion. We conduct an ablation study on Adv-BMT plus the forward refinement method, and results indicate a slightly better performance in average rewards, and average cost compared to Adv-BMT.
From closed-loop learning shown in table 4, we can see that agent learned with Adv-BMT environments has a significantly lower driving trajectory costs and higher rewards compared to baseline methods. With last step costs, collision rates, and driving completion rate remain consistent with the baseline. Comparing across tables with open-loop evaluations, we can see improvements over all metrics including rewards, completion rates, collision rates, and trajectory costs. This validates the value of diverse safety-critical scenarios which is provided in the closed-loop setting.
5 Conclusion and Limitations
Conclusion
This work introduces Adv-BMT, a novel safety-critical scenario generation framework that leverages BMT model to bring diverse and realistic adversarial collisions into real-world traffic scenarios. Adv-BMT consists of three key components: (1) collision initialization, which samples a potential collision outcome as the last frame; (2) reverse prediction, which generates the adversarial pre-collision trajectories; (3) rule-based filtering, which ensures adversarial agents’ motions remain realistic and feasible. We evaluate Adv-BMT across multiple benchmarks and experiments and demonstrate that training with generated adversarial scenarios improves AD agent performance in both held-out log-replay and collision test sets with a clear margin.
Limitations
Following limiations of our current work will be potentially addressed in the future: (1) we only consider adversarial behaviors among vehicle agents, but not among other agent types like pedestrains and bicycles. (2) we only evaluate our adversarial scenarios in RL policies, but adversarial scenarios can also be leveraged in human-in-the-loop learning.

### v2

@@ 3.3 Adv-BMT for Safety-critical Interaction
While existing works select a convenient neighbor agent and utilize its initial information, Adv-BMT inserts new agents (ADV) with diverse collision initializations for different opponent interactions. We visualize an example result in 5. Formally, we define the collision state as to include the position, time, velocity, and heading at the collision step. The collision time can be varied, sampled from the first second to the last time step of the ego trajectory length. Similarly, ADV’s collision headings are randomly sampled. ADV’s collision position can be calculated from collision heading and the ego vehicle’s collision position. Last but not least, collision velocity is calculated from sampling a offset from ego vehicle’s speed at collision step.
Multi-agent Adversarial Interaction.
A key advantage of Adv-BMT lies in its ability to generate realistic and flexible multi-agent interactions surrounding the adversarial agent (ADV). Starting from sampled collision initializations, the ADV is added into the scene together with existing traffic participants. BMT then performs reverse-time prediction to reconstruct plausible interaction histories that lead to the designated collision state. To balance realism, controllability, and scene reactivity, Adv-BMT provides several generation modes with different levels of multi-agent coupling:
  1. 1.
    Adv-BMT with replayed traffic agents: In this mode, traffic agents are teacher-forced to follow their original recorded trajectories, while only the adversarial agent’s history is predicted in reverse time. This preserves the original traffic flow and ensures high scene fidelity. Teacher-forcing prevents unintended deviations (e.g., traffic agents switching lanes or diverging from the original trajectory) that may occur if the model predicts all agents in the closed-loop manner. As a result, BMT reconstructs the adversarial agent’s trajectory to fit into the original scenario.
  2. 2.
    Adv-BMT with closed-loop reverse prediction: This mode removes teacher-forcing and jointly predicts the trajectories of all agents, including both traffic participants and the adversarial agent, in a single reverse prediction pass. By allowing all agents to evolve backward in time simultaneously, the scene develops coherently as a fully interactive scenario. This setting enables rich multi-agent dynamics and can produce fully synthesized safety-critical scenarios.
  3. 3.
    Adv-BMT with forward refinement: This hybrid mode first applies teacher-forced reverse prediction to generate the adversarial agent, then runs an additional forward prediction pass to enable traffic agents to react to its presence. This introduces interaction and reactivity without discarding the structure of the original recorded scene, offering a middle ground between strict replay and fully closed-loop generation.
These complementary modes provide a controllable trade-off between scene realism, interactivity, and controllability, enabling users to select the most suitable configuration for different evaluation objectives.
Rule-based Rejection Sampling.
@@
We design Adv-BMT to have diverse initializations, which do not guarantee the realism of the collision outcomes. To address this issue, we implement a rule-based filtering rejection mechanism for ADV candidates. Specifically, we first measure the driving distance and average speed of ADV candidates; if it moves too short and mostly wanders at the designated position waiting for the ego vehicle, then it is considered invalid. Meanwhile, we check the max curvature (the rate of change for heading): given a candidate ADV trajectory, we compute the curvature constraint using: $\kappa_{t}={\Delta\theta_{t}}/{\Delta s_{t}},$ where $\Delta\theta_{t}$ is the absolute heading change between time steps, and $\Delta s_{t}$ is the displacement between consecutive positions. Adv-BMT rejects a prediction if $\max_{t}(\kappa_{t})>\kappa_{\text{threshold}}$, where $\kappa_{\text{threshold}}=0.8$ is a predefined curvature limit that we found useful. We enforce a curvature constraint to ensure that the trajectory remains within the predefined threshold, rejecting ADV candidates that exhibit unrealistically sharp turning behavior. With our straightforward rule-based rejection sampling mechanism, we are able to maintain realistic collision events between the ADV and the ego vehicle.
@@ 4 Experiments
We first assess the BMT model on its ability to generate realistic and diverse traffic behaviors in Section 4.1. We then evaluate quality of Adv-BMT scenarios compared to three baseline methods in Section 4.2. Furthermore, in Section 4.3 we evaluate the utility for the downstream learning task, and assess whether training a reinforcement learning (RL) planner in Adv-BMT scenarios lead to improved performance and robustness compared to log-replay traffic flows.
Dataset.
All experiments use driving data from the Waymo Open Motion Dataset (WOMD) ettinger2021largescaleinteractivemotion with formats managed by ScenarioNet li2023scenarionetopensourceplatformlargescale . WOMD contains 10Hz scenarios, each with 1 second of history and 8 seconds of future trajectories. Each scenario includes up to 128 traffic agents including vehicles, cyclists, pedestrians along with high-definition maps. To reduce computational cost, we downsample each scenario to 2Hz, yielding 19 prediction steps. We randomly select 500 scenes for both open-loop evaluation and RL training. We use 6 prediction modes for each scene during open-loop evaluations in Section 4.1 and Section 4.2.
Metrics.
To evaluate the realism of BMT predictions, we adopt standard open-loop prediction accuracy metrics: Scenario Final Displacement Error (SFDE) and Scenario Average Displacement Error (SADE). For each, we report results on both the best prediction mode and the average over all six prediction modes. We also measure the average number of agent collisions to capture interaction realism. To assess the diversity of generated trajectories, we report Final Displacement Diversity (FDD), Starting Displacement Diversity (SDD), and Average Displacement Diversity (ADD), which quantify the spread of predicted positions across all prediction modes. We also compute the Jensen–Shannon Divergence (JSD) values over velocity, acceleration, and time-to-collision (TTC) distributions between generated and ground-truth trajectories. JSD measures the similarity between two probability distributions, with lower values indicating closer alignment between generated and real-world behaviors.
4.1 Evaluation of BMT Predictions
Table 1: We evaluate BMT under three types of prediction tasks in a closed-loop setting for all traffic agents, without replaying any agents. In the bidirectional mode, forward prediction is performed based on the reverse prediction results. The number of prediction modes is set to six.
(a) Realism of BMT predictions.
Method $_{\text{avg}} SFDE$$_{\text{min}}$ $_{\text{avg}} SADE$$_{\text{min}}$ $_{\text{min}} VehColl$$_{\text{avg}}$ $_{\text{velocity}} JSD$$_{\text{TTC}}$
Reverse 3.92 2.70 5.53 5.11 0.03 0.05 0.17 0.02
Forward 3.52 2.35 2.39 1.98 0.03 0.05 0.23 0.30
Reverse + Forward 3.29 2.01 2.38 1.74 0.03 0.06 0.22 0.35
(b) Diversity of BMT predictions.
Method SDD FDD ADD
Reverse 8.35 - 3.13
Forward - 10.78 4.40
Reverse + Forward 8.49 12.78 7.00
@@ 4.1 Evaluation of BMT Predictions
Results from Table 1(a) indicates that BMT performs realistic scenario generations in forward, backward, and bidirectional predictions. Reverse predictions perform slightly worse than forward predictions in realism metrics. Compared to single-pass prediction, bidirectional prediction exhibit higher prediction accuracy. Despite these differences, the overall collision rate remains reasonable and comparable for all prediction tasks, which indicates that BMT effectively generate realistic traffic interactions. The results in Table 1(b) show that bidirectional prediction enhances the diversity of generated agent behaviors across all displacement diversity metrics. Forward prediction exhibits higher FDD and ADD scores than reverse prediction. This indicates that predicting future motions encourages more explorations for variant directions.
4.2 Evaluation of Adv-BMT Scenarios
In Table 2, we additionally use the adversarial attack success rate (i.e., the collision rate between adversary and the ego), ADV-Traffic Collision Rate (i.e.,collision rates between adversary agents and traffic agents), and the average Agent Collision (i.e., average traffic agent collision rate) to indicate interaction realism. Two settings of Adv-BMT outperform baselines in both realism and diversity. This validates Adv-BMT’s design of collision initialization + reverse predictions, which couldn’t be achieved by other methods. While Adv-BMT generates highly adversarial behaviors, it also preserves diversity of traffic interactions compared to baselines. FDD and SDD metrics suggest that baselines generate nearly the same adversarial trajectories on the given scene. The JSD metrics suggest Adv-BMT outperforms in realism metrics compared to baselines. BMT model is able to achieve realistic motion predictions indicted by Waymo Open Sim Agent Challenge metrics montali2023waymoopensimagents , which we add in Table 6 in our appendix. Furthermore, the results for Adv-BMT with filtering indicate that our rule-based filter does not harm diversity and at the same time enhances the realism metrics. Table 3 demonstrates the generation speed comparisons among all methods. Our evaluations validate Adv-BMT as an efficient framework for realistic, diverse, and safety-critical generation.
Table 2: Diversity and realism of generated adversarial behaviors. We compare Adv-BMT scenarios with three baseline methods for safety-critical interactions, namely STRIVE rempe2022generatingusefulaccidentpronedriving , CAT zhang2023catclosedloopadversarialtraining , and SEAL stoler2025sealsafeautonomousdriving . To be consistent with baseline methods in this evaluation only, we force Adv-BMT to choose nearest neighboring agent and modify its behavior as the predicted adversary.
@@ 4.2 Evaluation of Adv-BMT Scenarios
CAT 0.00 0.00 0.13 0.16 0.12 0.48 0.10 0.3
SEAL 0.00 0.00 0.18 0.25 0.17 0.55 0.12 0.32
STRIVE 0.01 0.00 0.11 0.22 0.12 0.38 0.09 0.32
BMT_TF 1.22 0.49 0.10 0.17 0.13 1.00 0.13 0.22
BMT_All 1.51 0.63 0.10 0.17 0.08 1.00 0.13 0.19
BMT_TF + Filter 2.32 1.00 0.08 0.12 0.09 1.00 0.13 0.12
BMT_All + Filter 1.97 0.83 0.13 0.14 0.08 1.00 0.12 0.09
Visualization.
We simulate Adv-BMT scenarios in MetaDrive li2022metadrivecomposingdiversedriving , rendering in both bird’s-eye view and third-person perspective. In Fig. 5, Fig. 4, and Fig. 9, we show diverse Adv-BMT adversarial behaviors in the collision directions on the same ego agent in several real-world traffic flows from WOMD. Generated adversarial agents follow traffic rules and maintain realistic driving patterns. Adv-BMT supports diverse adversarial agent types including vehicles, pedestrians, and cyclists, as shown in Fig. 6. Adv-BMT makes full use of each driving log, which makes Adv-BMT suitable for AD testing and adversarial training. A visual comparison between Adv-BMT and baseline methods is shown in Fig. 7. When a baseline fails to select an existing traffic agent that is convenient for a safety-critical attack, Adv-BMT is able to imagine a new agent at an appropriate position to perform adversarial attacks on the ego vehicle.
Table 3: Generation speeds across methods. Among the four methods, CAT achieves the fastest generation speed, followed by Adv-BMT with a slightly lower speed. SEAL and STRIVE are comparatively slower.
Method CAT SEAL STRIVE Adv-BMT
500-avg (seconds) 0.80 2.36 9.53 1.02
@@ 4.3 Adversarial Learning
To validate the value of Adv-BMT scenarios in downstream autonomous driving (AD) tasks, we train a reinforcement learning (RL) agent within augmented scenarios containing collision interactions generated by Adv-BMT and baseline methods. To determine the quality of the augmented training scenarios, we measure both the driving performance and the safety performance of the learned AD agent compared to learning on the original training set. In our experiment, we conduct two sets of training: (1) open-loop RL, where the agent is trained on a fixed training set with adversarial scenarios generated based on ground-truth ego trajectories; and (2) closed-loop RL, where an adaptive adversarial agent attacks the current ego agent based on its recent rollout trajectory records. The adaptive adversarial agent’s motion is generated by Adv-BMT or a baseline method. Results are shown in Table 4 and Table 5.
Setting.
The training set contains 500 real-world scenarios randomly selected from the WOMD training set. We train a Twin Delayed DDPG (TD3) agent for 1 million steps using 8 random seeds to ensure robustness in MetaDrive li2022metadrivecomposingdiversedriving (hyperparameters listed in the appendix). We measure the average reward, average step cost, average route completion rate (Compl.), and average episode cost (cost sum) for driving performance measurement. To evaluate the impact of adversarial training using Adv-BMT-generated scenarios, we assess policy performance across two distinct validation environments: (1) 100 Waymo validation environments, which consist of unmodified real-world driving scenarios from WOMD validation set, and (2) 100 Adv-BMT environments, which is the augmented collision scenarios from the 100 validation scenes.
Table 4: Open-loop RL agent evaluation. Each WOMD scenario is augmented with one collision scenario, so that the ratio between real-world and safety-critical scenes is 1:1. For each method, we augment the original dataset into a new training set.
@@
Training Scenarios Reward $\uparrow$ Cost $\downarrow$ Completion $\uparrow$ Collision $\downarrow$ Cost Sum $\downarrow$
@@
Adv-BMT (Refined) 33.22 $\pm$ 1.83 0.36 $\pm$ 0.03 0.74 $\pm$ 0.03 0.12 $\pm$ 0.02 1.39 $\pm$ 0.22
(b) Evaluation in the Adv-BMT validation environments
Training Scenario Reward $\uparrow$ Cost $\downarrow$ Completion $\uparrow$ Collision $\downarrow$ Cost Sum $\downarrow$
@@
Adv-BMT (Refined) 39.55 $\pm$ 2.94 0.59 $\pm$ 0.04 0.65 $\pm$ 0.02 0.27 $\pm$ 0.04 2.74 $\pm$ 0.54
Table 5: Closed-loop RL agent evaluation. We use the same augmented training set as in the open-loop experiment. For the adaptive adversarial learning experiment, we implement an adaptive generator for Adv-BMT and CAT zhang2023catclosedloopadversarialtraining . We discard the other baseline methods due to their low generation speeds.
@@
Generator Reward $\uparrow$ Cost $\downarrow$ Completion $\uparrow$ Collision $\downarrow$ Cost Sum $\downarrow$
@@
Generator Reward $\uparrow$ Cost $\downarrow$ Completion $\uparrow$ Collision $\downarrow$ Cost Sum $\downarrow$
@@
Analysis.
Table 4 reports open-loop evaluations comparing RL policies trained on Adv-BMT–generated scenarios against policies trained on baseline scenario sets. On the WOMD original validation environments, Adv-BMT–trained agents outperform all baselines across metrics: episode cost decreases by 10% and collision rate by 8%, while reward and route-completion remain comparable to the best baseline. Adding the optional forward refinement (Sec. 3.3) yields additional gains—reward increases by 6%, episode cost decreases by 5%, and route completion improves slightly. On safety-critical test environments, Adv-BMT further reduces episode cost by 17% and collision rate by 14% relative to the best baseline. With forward refinement, we observe further improvements in driving capability, reflected in higher reward and completion rates. These results indicate that training on Adv-BMT scenarios exposes agents to a broader distribution of high-risk interactions, enabling more robust, safety-aware policies. The extra gains from forward refinement underscore the importance of modeling reactive traffic around emerging incidents, which better prepares agents for cascading hazards. Motivated by these findings, we next evaluate all methods in a closed-loop RL setting.
Table 5 reports closed-loop RL results under reactive adversarial traffic generators. Relative to the corresponding open-loop policy, the closed-loop policy improves across all metrics in both real-world and safety-critical collision environments, indicating a substantially safer and more robust AD policy. Compared with the strongest closed-loop baseline (CAT), our method achieves higher reward, lower collision rate, and a 38% reduction in episode cost. These gains suggest that training with Adv-BMT scenarios—featuring diverse adversarial behaviors and realistic, traffic-consistent reactions—better train the agent to anticipate and mitigate high-risk interactions. Learning in closed-loop further improves performance; by training the agent to respond online to reactive opponents, it becomes better at anticipating dangerous situations, and staying safe in unfamiliar or rapidly changing traffic conditions. The results also highlight the importance of the Adv-BMT design, which enables the generation of realistic and flexible multi-agent interactions with the adversarial agent.
Figure 6: Adv-BMT can also generate scenarios of pedestrian and cyclist adversarial agents.
Figure 7: Baseline method comparisons. We present several examples where Adv-BMT successfully generates collision interactions in scenarios where all baseline methods fail. Unlike the baselines, Adv-BMT does not rely on the proximity of neighboring agents to the ego vehicle, enabling more flexible and diverse adversarial attack strategies.
5 Conclusion
We introduced Adv-BMT, a novel framework for generating diverse safety-critical driving scenarios with realistic traffic interactions. Built on a Bidirectional Motion Transformer with the ability to perform bidirectional motion prediction tasks, Adv-BMT is able to design a new adversarial agent for each real-world scenario. A key advantage of Adv-BMT lies in its ability to generate realistic and flexible multi-agent interactions surrounding the adversarial agent. Unlike previous frameworks that select an existing neighboring vehicle and modify the corresponding trajectory for an adversarial attack, Adv-BMT initializes candidate collisions for a new adversarial agent, reconstructs multi-agent trajectories via inverse prediction, and optionally applies forward refinement for reactive, traffic-consistent interactions. Adv-BMT is able to balance realism, controllability, and scene reactivity. Our evaluations validate the quality of Adv-BMT scenarios in terms of interaction realism and diversity; furthermore, we demonstrate that learning within the Adv-BMT traffic flow improves AD agents’ safety performance by a clear margin.
Limitations.
Adv-BMT relies on long token sequences to represent multi-agent traffic scenes, which results in high memory and computational demands. Our empirical study focuses on a single downstream task—reinforcement learning for an autonomous driving agent—so broader applicability remains to be validated.
Acknowledgment: This work was supported by the NSF Grants CNS-2235012, IIS-2339769, and CCF-2344955. ZP was supported by the Amazon Fellowship via UCLA Science Hub.